OpenStudy (anonymous):

2x + y = 7 5x + y = 9 Solve the system of equations. @Kemist can anybody help?

OpenStudy (igreen):

(0.667), (5,667) https://www.desmos.com/calculator

OpenStudy (muzzack):

hey @iGreen please dont tell them like that explain why

OpenStudy (anonymous):

@stocksro https://www.desmos.com/calculator works out alot for your system of equations.

OpenStudy (muzzack):

so first we start like this \(2x + y = 7\) \(5x + y = 9\) --------- then we first make the first equation isolate y --------- \(y=-2x+7\) --------- then replace the 2nd equation with the first --------- \(5x+1(-2+7)=9\)

OpenStudy (muzzack):

then we get --------- 5x-2+7=9

OpenStudy (muzzack):

now what we do is solve for x -------- 5x+5=9 ------- 5x=4 ------- x=4/5